# A New Census of the Universe’s Entropy

### v1

@@ 8.4 Maximal Entropy of Super-massive Primordial Black Holes
Fig. 4 shows the maximal entropy, compatible with current constraints on the PBH abundance at a given mass, for a variety of lognormal mass function widths $\sigma$ (see Eq. (8.1)). The right panel zooms in on the very large mass regime. While Fig. 4 assumes that no constraints exist past the CMB distorsion limit, Fig. 5 assumes that the constraints are constant past the CMB distorsion limit (purple line in the previous figures).
@@ 8.5 Planck-scale relics
We thus conclude that if the dark matter consists of Planck-scale relic the associated entropy is extremely small, with a comparable magnitude to the entropy hosted in stellar remnants like neutron stars. Note that if the relics are charged, the entropy will be increasingly smaller as the event horizon of charged black holes of a given mass is smaller than that of a non-charged black hole.
@@ 9 Discussion and Conclusions
Figure 6: A summary of the total entropy in the universe associated with the various components described in this study.
@@
    within the class of stellar objects and remnants, by far the lion share of the entropy is in main-sequence light, solar-mass stars, with heavier main-sequence stars and white dwarfs contributing roughly three orders of magnitude less, and neutron stars roughly eight orders of mangitude less;
@@
    for the first time, we computed the entropy contribution from cosmic ray hadrons and leptons; the largest source of uncertainty there is to extrapolate from local measurements to the universe as a whole, amounting to roughly, we estimated, two orders of magnitude; we found that cosmic ray protons contribute two orders of mangitude more than electrons, and their contribution to the global entropy budget could be as large as 10% of that of baryons in bound objects such as stars;
@@
    a major novelty of our study is a comprehensive review of the possible entropy stashed in the cosmological dark matter sector. We have shown that if the dark matter sector is small, the entropy associated with it is, at most, as large as that of the CMB; however, we pointed out two important caveats: first, late time clustering could enhance the dark matter entropy by up to six orders of magnitude; second, a large dark sector could potentially make the dark matter sector entropy significantly larger than that of any other component in the universe, and almost as large than that of the cosmic event horizon, with deep and dramatic potential implications for the thermodynamics of the universe and for the GSL;
@@
    we re-assessed in detail the contribution of stellar, intermediate-mass, and supermassive black holes, in light of new results on the black hole mass functions in the respective mass ranges. As previously thought, the Bekenstein-Hawking entropy associated with supermassive black holes dominate the (known) entropy budget of the universe, with a total entropy that, we estimate, exceeds $10^{100}\ k_{B}$; the entropy associated with IMBH is about 4 orders of magnitude less, and that of stellar-mass black hole 5;
@@
    for the first time, we computed the maximal amount of entropy that could possible be associated with primordial black holes, i.e. black holes of non-stellar origin. For masses up to a few thousand solar masses, where constraints on the PBH abundance are robust, we found that, depending on the PBH mass function, the maximal entropy in PBH tops at slightly less than $10^{100}\ k_{B}$, and can thus be as large as that in SMBHs. However, a much larger amount of entropy could be associated with supermassive primordial black holes. We find that, depending on the assumed constraints, supermassive PBH can be associated with as much as $10^{118}\ k_{B}$, almost a billion times more than that of ordinary SMBH, and just a few orders of mangitude below the entropy of the CEH;
@@
We show in our plot in fig. 6 the various entropy components we surveyed here, in order to give a final visual summary of our study’s main quantitative findings.
@@ Acknowledgments
This work is partly supported by the U.S. Department of Energy grant number de-sc0010107 (SP). SP thanks Nicholas Profumo for help with references typesetting.

### v2

@@ 8.5 Maximal Entropy of Super-massive Primordial Black Holes
To provide intuition for the proximity of supermassive primordial black hole (SMPBH) entropy to that of the cosmic event horizon, we present here a simplified estimate.
Assume that there exist $N$ SMPBH of mass $M$ such that they contribute a fraction $f=NM/M_{H_{0}}$ of the current total mass density
$M_{H_{0}}\sim\rho_{0}\frac{4\pi}{3}H_{0}^{-3}$
in the universe, where hereafter for simplicity we adopt Planck units ($G=\hbar=c=k_{B}=1$). Note that since $\rho_{0}=\frac{3H_{0}^{2}}{8\pi},$ we have that $M_{H_{0}}=H_{0}^{-1}/2$. For comparable dark matter and dark energy densities, the entropy associated with the SMPBHs, $S_{\mathrm{SMPBH}}=N\cdot 4\pi M^{2}=4\pi fM_{H_{0}}M$, will thus be a fraction of order
$\frac{S_{\rm SMPBH}}{S_{\rm CEH}}\sim f\frac{M}{M_{H_{0}}}$ (8.9)
of the entropy associated with the CEH. The left panel of fig. 2 suggests that for the largest-possible SMPBH, at the intersection of the “incredulity limit” and the CMB limits, $M/M_{H_{0}}\lesssim 10^{-2}$ and $f\lesssim 10^{-2}$, thus predicting an upper limit to the ratio $S_{\rm SMPBH}/S_{\rm CEH}\lesssim 10^{-4}$. Our detailed analysis reflects quite closely this order-of-magnitude estimate [^1] We are grateful and indebted to the Anonymous Referee for suggesting, and indeed for explicitly working out, this estimate.. Notice that $S_{\rm SMPBH}/S_{\rm CEH}\sim 10^{-4}$, for $M/M_{H_{0}}\lesssim 10^{-2}$ and $f\lesssim 10^{-2}$, would correspond to $N\sim 1$ SMPBHs.
Generally, for monochromatic black holes of mass $M$ constituting a fraction $f$ of the dark matter, we find that the corresponding entropy is
$S(f,M)\simeq 5\times 10^{114}\ k_{B}\cdot f\cdot\left(\frac{M}{10^{50}\ {\rm g}}\right).$ (8.10)
In turn, this implies, for instance for the choice of a “constant” CMB constraint yielding a maximal mass at the incredulity limit of $M=6\times 10^{47}{\rm g}$ and $f\simeq 3\times 10^{-9}$, an entropy of $\sim 10^{104}\ k_{B}$; in the absence of CMB constraints (rightmost corner of the grey regions in the left panel of fig. 2-3), SMPBH can be as massive as $6\times 10^{54}\ {\rm g}$ with $f\simeq 3\times 10^{-2}$, harboring an entropy as large as $S\sim 10^{118}\ k_{B}$. This latter figure is consistent with the estimate above for $N\sim 1$ SMPBH and with our numerical findings in the previous section.
Fig. 4 shows the maximal entropy, compatible with current constraints on the PBH abundance at a given mass, for a variety of lognormal mass function widths $\sigma$ (see Eq. (8.4)). The right panel zooms in on the very large mass regime. While Fig. 4 assumes that no constraints exist past the CMB distortion limit, Fig. 5 assumes that the constraints are constant past the CMB distortion limit (purple line in the previous figures).
@@ 8.6 Planck-scale relics
We thus conclude that if the dark matter consists of Planck-scale relics the associated entropy is extremely small, with a comparable magnitude to the entropy hosted in stellar remnants like neutron stars. Note that if the relics are charged, the entropy will be increasingly smaller as the event horizon of charged black holes of a given mass is smaller than that of a non-charged black hole.
@@ 9 Discussion and Conclusions
The transition from inflation to the hot Big Bang phase, known as reheating, represents one of the most dramatic entropy-generating epochs in cosmic history. During inflation, the universe exists in a quasi-de Sitter state dominated by the potential energy of the inflaton field $\phi$, characterized by a nearly constant Hubble parameter. This state has remarkably low entropy density, primarily associated with the de Sitter horizon, as the matter content is restricted to the highly homogeneous inflaton field with negligible thermal entropy. As inflation ends, the inflaton begins oscillating around the minimum of its potential, triggering several entropy-generating mechanisms:
  • Perturbative decays: The inflaton field $\phi$ decays into Standard Model particles with a rate $\Gamma_{\phi}$, gradually transferring its energy to radiation with $\rho_{r}\propto a^{-4}$ and generating entropy.
  • Parametric resonance: In many inflation models, the early oscillation phase triggers exponential particle production through parametric resonance, often called ”preheating”. This non-perturbative process rapidly converts inflaton energy into out-of-equilibrium particles within specific momentum bands.
  • Thermalization: The produced particles subsequently interact and scatter, approaching thermal equilibrium and maximizing the coarse-grained entropy within the available energy constraints.
Quantitatively, the entropy density increases by a factor of approximately $e^{3N}$ during reheating, where $N\approx 50$$60$ is the number of e-folds during inflation.
The entropy production during reheating establishes the initial conditions for the conventional hot Big Bang phase, with total entropy now primarily carried by relativistic particles. The massive entropy generation during reheating thus sets the stage for all subsequent thermodynamic evolution of the universe, establishing the high-entropy background from which complex structures later emerge through gravitational collapse.
@@
[…]
@@
    within the class of stellar objects and remnants, by far the lion share of the entropy is in main-sequence light, solar-mass stars, with heavier main-sequence stars and white dwarfs contributing roughly three orders of magnitude less, and neutron stars roughly eight orders of magnitude less;
@@
    for the first time, we computed the entropy contribution from cosmic ray hadrons and leptons; the largest source of uncertainty there is to extrapolate from local measurements to the universe as a whole, amounting to roughly, we estimated, two orders of magnitude; we found that cosmic ray protons contribute two orders of magnitude more than electrons, and their contribution to the global entropy budget could be as large as 10% of that of baryons in bound objects such as stars;
@@
    a major novelty of our study is a comprehensive review of the possible entropy stashed in the cosmological dark matter sector. We have shown that if the dark matter sector is small, the entropy associated with it is, at most, as large as that of the CMB; however, we pointed out two important caveats: first, late time clustering could enhance the dark matter entropy by up to six orders of magnitude; second, a large dark sector could potentially make the dark matter sector entropy significantly larger than that of any other component in the universe, and almost as large as that of the cosmic event horizon, with deep and dramatic potential implications for the thermodynamics of the universe and for the GSL;
@@
    we re-assessed in detail the contribution of stellar, intermediate-mass, and supermassive black holes, in light of new results on the black hole mass functions in the respective mass ranges. As previously thought, the Bekenstein-Hawking entropy associated with supermassive black holes dominate the (known) entropy budget of the universe, with a total entropy that, we estimate, exceeds $10^{100}\ k_{B}$; the entropy associated with IMBHs and stellar-mass black holes is about 4 orders of magnitude less than that of supermassive black holes;
@@
    for the first time, we computed the maximal amount of entropy that could possible be associated with primordial black holes, i.e. black holes of non-stellar origin. For masses up to a few thousand solar masses, where constraints on the PBH abundance are robust, we found that, depending on the PBH mass function, the maximal entropy in PBH tops at slightly less than $10^{100}\ k_{B}$, and can thus be as large as that in SMBHs. However, a much larger amount of entropy could be associated with supermassive primordial black holes. We find that, depending on the assumed constraints, supermassive PBH can be associated with as much as $10^{118}\ k_{B}$, almost a billion times more than that of ordinary SMBH, and just a few orders of magnitude below the entropy of the CEH;
@@
We show in our plot in fig. 6 the various entropy components we surveyed here, in order to give a final visual summary of our study’s principal quantitative findings. To complement the graphical presentation in fig. 6, we provide in Table LABEL:tab:tab1 a numerical summary of our entropy estimates for all components considered in this study, including associated uncertainties where available. The fourth and fifth column present a side-by-side comparison of our updated values with those reported in previous studies—primarily Ref. [11]—highlighting the effects of updated cosmological parameters, new astrophysical data, and methodological improvements.
[…]
Component $\log_{10}(S/k_{B})$ Unc. (dex) Ref. [11] Unc. (dex)
CEH 122.48 0.05 122.10 0.05
SMBH 103.25 0.19 102.50 0.30
Stellar BHs 101.50 0.09 101.20 0.05
CMB 89.72 0.03 89.69 0.03
[…]
Stellar matter 81.98 0.20 81.50 0.10
Diffuse baryons 83.15 0.34 83.00 0.30
Cosmic Rays 78.50 1.00 * *
Thermonuclear $\nu$’s 74.50 1.00 * *
Dark Matter (thermal) 88.30 0.43 85.00 1.00
Dark Sector (max) 128.00 1.00 * *
@@ Acknowledgments
This work is partly supported by the U.S. Department of Energy grant number de-sc0010107 (SP). We thank Batoul Banihashemi and Edgar Shaghoulian for helpful discussions. SP thanks Nicholas Profumo for help with references typesetting.